40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 24 01:16:30 2015, Jenne, Update, ASC, Broken Xass? 
    Reply  Sat Apr 25 03:32:45 2015, Koji, Update, ASC, Broken Xass? 
       Reply  Sun Apr 26 01:10:18 2015, rana, Update, ASC, unBroken Xass? xasstune_150426.pngxassnoise.pdf
          Reply  Sun Apr 26 15:05:35 2015, Jenne, Update, ASC, unBroken Xass? 
Message ID: 11248     Entry time: Sat Apr 25 03:32:45 2015     In reply to: 11242     Reply to this: 11253
Author: Koji 
Type: Update 
Category: ASC 
Subject: Broken Xass? 

I spent a day to fix the XARM ASS, but no real result. If the input of the 6th DOF servo is turned off, the other error signals are happy
to be squished to around their zeros. So this gives us some sort of alignment control. But obviously a particular combination of the
misalignment is left uncontrolled.

This 6th DOF uses BS to minimize the dither in ITMX yaw. I tired to use the other actuators but failed to have linear coupling between
the actuator and the sensor.


During the investigation, I compared TRX/TRY power spectra. TRX had a bump at 30Hz. Further investigation revealed that the POX/POY
had a big bump in the error signals. The POX/POY error signals between 10-100Hz were coherent. This means that this is coming from
the frequency noise stabilized with the MC. (Is this frequency noise level reasonable?)

The mysterious discovery was that the bump in the transmission exist only in TRX. How did the residual frequency noise cause
the intensity noise of the transmission? One way is the PDH offset.

Anyway, Rana pointed out that IMC WFS QPDs had large spot offsets. Rana went to the AS table and fixed the WFS spot centering.
This actually removed the bump in TRX although we still don't know the mechanism of this coupling.

The bump at 30Hz was removed. However, the ASS issue still remains.

ELOG V3.1.3-